OpenStudy (anonymous):

Solve the equation. (Round your answer to three decimal places if necessary.) x2/5+5x1/5=-6

OpenStudy (anonymous):

this is a quadratic equation in \[x^{\frac{1}{5}}\]

OpenStudy (anonymous):

put say \[z=x^{\frac{1}{5}}\] and get \[z^2+5z=-6\] \[z^2+5z+6=1\] \[(z+3)(z+2)=0\] \[z=-3, z=-2\]

OpenStudy (anonymous):

thens solve \[x^{\frac{1}{5}}=-3\] \[x=(-3)^5\] etc